OpenStudy (harleygirl18):

Pretty Please help with a cherry on top, Don't understand this at all will fan and medal. What is the side length of a square with an area of 225 cm2? Use an appropriate technique to solve. A. 112.5 cm B. 15 cm C. 56.25 cm D. 45 cm

zepdrix (zepdrix):

Area of a rectangle is given by length time width: \(\large\rm A=\ell\cdot w\) If we have a square, then the length and width are the same: \(\large\rm A=\ell\cdot\ell\) We can write the length times length as length squared: \(\large\rm A=\ell^2\) They gave us the value of the area, so we can plug it in, \(\large\rm 225=\ell^2\)

zepdrix (zepdrix):

So to solve for the side length l, we need to "undo" the squaring on it.

zepdrix (zepdrix):

Do you remember how to undo "squaring"? Hint: we undo `multiplication` with `division` we undo `addition` with `subtraction` we undo `square` with ?

OpenStudy (harleygirl18):

i don't know how to do squaring

zepdrix (zepdrix):

Anyway, you undo `squaring` with `square root`. Applying square root to each side gives us,\[\large\rm \sqrt{225}=\sqrt{\ell^2}\]The square root and square undo one another, giving us our side length.\[\large\rm \sqrt{225}=\ell\] And then you'll have to uh, simply a little bit from there.